### v2

@@ 8.1. Extension graphs in codings of IET
A shift space $X$ over $\mathcal{A}$ is the natural coding of a regular interval exchange transformation with the pair of orders $\binom{\leq_{1}}{\leq_{2}}$ if and only if it is minimal, $\mathcal{A}\subset\mathcal{L}(X)$ and it satisfies the following conditions for every $w\in\mathcal{L}(X)$:
@@
This result can be reformulated as follows. Let $\leq^{L}$ and $\leq^{R}$ be two total orders on $\mathcal{A}$. A factor $w$ of a shift space $X\subset A^{\mathbb{Z}}$ is said to be planar for $(\leq^{L},\leq^{R})$ if item 1 is satisfied for $\binom{\leq^{R}}{\leq^{L}}$. This implies that, placing the left and right vertices of $\mathcal{E}_{X}(w)$ on parallel lines and ordering them respectively by $\leq^{L}$ and $\leq^{R}$, the edges of $\mathcal{E}_{X}(w)$ may be drawn as straight noncrossing segments, resulting in a planar graph. In particular, $\mathcal{E}_{X}(w)$ is acyclic.
@@
    Any long enough left special factor $w$ is such that $E^{L}_{X}(w)$ is equal to $\{a,b\}$, where $a,b$ are consecutive for $\leq^{L}$. Furthermore, for every two $\leq^{L}$-consecutive letters $a,b$ and for all $n\in\mathbb{N}$, there is a (unique) left special factor $w$ of length $n$ such that $\{a,b\}\subset E_{X}^{L}(w)$.
@@
    The same holds on the right, i.e., any long enough right special factor $w$ is such that $E^{R}_{X}(w)$ is equal to $\{a,b\}$, where $a,b$ are consecutive for $\leq^{R}$. Furthermore, for every two $\leq^{R}$-consecutive letters $a,b$ and for all $n\in\mathbb{N}$, there is a (unique) right special factor $w$ of length $n$ such that $\{a,b\}\subset E_{X}^{R}(w)$.
@@ 8.2. S-adic characterization of regular interval exchange transformations
From Theorem 63, we know that the natural coding of a regular interval exchange transformation is a minimal dendric shift. By Theorem 1, any of its $\mathcal{S}$-adic representation made of return morphisms labels a path in $\mathcal{G}^{L}(\mathcal{S})$ and in $\mathcal{G}^{R}(\mathcal{S})$. In this section, we characterize those labeled paths.
@@
Let $a,b$ be two distinct letters in $\mathcal{A}$. Let $\mathcal{S}$ be a set of dendric return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$. The graph $\mathcal{G}_{IET}(\mathcal{S})$ is such that its vertices are the irreducible pairs of orders $(\leq^{L},\leq^{R})$ such that $a<^{L}b$, and there is an edge from $(\leq^{L},\leq^{R})$ to $(\preceq^{L},\preceq^{R})$ labeled by $\sigma\in\mathcal{S}$ if
@@
    $(G(\leq^{L}),\sigma,G(\preceq^{L}))$ is an edge of $\mathcal{G}^{L}(\mathcal{S})$;
@@
    $(G(\leq^{R}),\sigma,G(\preceq^{R}))$ is an edge of $\mathcal{G}^{R}(\mathcal{S})$;
@@
Let $\mathcal{S}$ be a family of dendric return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$ and let $X$ be a shift space having an $\mathcal{S}$-adic representation ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}$. Then $X$ is the coding of a regular interval exchange if and only if ${\boldsymbol{\sigma}}$ is primitive and labels an infinite path in the graph $\mathcal{G}_{IET}(\mathcal{S})$. Moreover, if this path starts in the pair $(\leq^{L},\leq^{R})$, then $X$ is the coding of a regular interval exchange for the orders $\binom{\leq^{R}}{\leq^{L}}$.
@@
Let us assume that $X$ is the coding of a regular interval exchange for the orders $\binom{\leq^{R}_{0}}{\leq^{L}_{0}}$, we can moreover assume that $a<^{L}_{0}b$, where $a$ and $b$ are the letters used to define $\mathcal{G}_{IET}(\mathcal{S})$. Then, $\sigma_{0}$ is $(\leq^{L}_{0},\leq^{R}_{0})$-planar, and by Lemma 67, there exist two orders $\leq^{L}_{1}$ and $\leq^{R}_{1}$ such that $a<^{L}_{1}b$ and $\sigma$ is planar preserving from $(\leq^{L}_{1},\leq^{R}_{1})$ to $(\leq^{L}_{0},\leq^{R}_{0})$ or from $((\leq^{L}_{1})^{*},(\leq^{R}_{1})^{*})$ to $(\leq^{L}_{0},\leq^{R}_{0})$.
As $X_{\boldsymbol{\sigma}}^{(1)}$ is dendric, by Corollary 70, it is the coding of a regular interval exchange for the orders $\binom{\leq^{R}_{1}}{\leq^{L}_{1}}$. In particular, the pair $(\leq^{L}_{1},\leq^{R}_{1})$ is irreducible. Moreover, $\sigma_{0}$ labels an edge from $G(\leq^{L}_{0})$ to $G(\leq^{L}_{1})$ in $\mathcal{G}^{L}(\mathcal{S})$ and from $G(\leq^{R}_{0})$ to $G(\leq^{R}_{1})$ in $\mathcal{G}^{R}(\mathcal{S})$. Therefore, $\sigma_{0}$ labels an edge from $(\leq^{L}_{0},\leq^{R}_{0})$ to $(\leq^{L}_{1},\leq^{R}_{1})$ in $\mathcal{G}_{IET}(\mathcal{S})$. We then iterate this reasoning to show that ${\boldsymbol{\sigma}}$ labels an infinite path in $\mathcal{G}_{IET}(\mathcal{S})$.
Assume now that ${\boldsymbol{\sigma}}$ is primitive and labels a path $((\leq^{L}_{n},\leq^{R}_{n}))_{n\geq 0}$ in $\mathcal{G}_{IET}(\mathcal{S})$. Since ${\boldsymbol{\sigma}}$ also labels an infinite path in $\mathcal{G}^{L}(\mathcal{S})$ and $\mathcal{G}^{R}(\mathcal{S})$, $X$ is minimal dendric. Let us prove that it is planar for $(\leq^{L}_{0},\leq^{R}_{0})$. The $\sigma_{0}$-initial factors of $X$ are planar for $(\leq^{L}_{0},\leq^{R}_{0})$ since $\sigma_{0}$ is $(\leq^{L}_{0},\leq^{R}_{0})$-planar. Let $u\in\mathcal{L}(X)$ be a non-$\sigma_{0}$-initial factor of $X$. There exist a unique $k>0$ and a unique $v\in\mathcal{L}(X_{\boldsymbol{\sigma}}^{(k)})$ $\sigma_{k}$-initial such that $u$ is an extended image of $v$ by $\sigma_{0}\dots\sigma_{k-1}$ by Proposition 8. However, $\sigma_{k}$ is $(\leq^{L}_{k},\leq^{R}_{k})$-planar thus $v$ is planar for these orders. Iterating Proposition 69 $k$ times allows us to conclude that $u$ is planar for $(\leq^{L}_{0},\leq^{R}_{0})$. As this is true for any $u\in\mathcal{L}(X)$, $X$ is the coding of a regular interval exchange for the orders $\binom{\leq^{R}_{0}}{\leq^{L}_{0}}$. ∎
@@ 9. Future work
Given a set $\mathcal{S}$ of dendric return morphisms, Theorem 1 characterizes the $\mathcal{S}$-adic shifts that are minimal and dendric. We thus directly obtain the following corollary.
@@
Let $\mathcal{S}$ be a set of dendric return morphisms defined on some alphabet $\mathcal{A}$ such that any minimal dendric shift over $\mathcal{A}$ admits an $\mathcal{S}$-adic representation. Then a shift space $X\subseteq\mathcal{A}^{\mathbb{Z}}$ is minimal and dendric if and only if it has a primitive $\mathcal{S}$-adic representation labeling paths in the graphs $\mathcal{G}^{L}(\mathcal{S})$ and $\mathcal{G}^{R}(\mathcal{S})$.
This corollary is however not completely satisfactory. Indeed it is in general difficult to find such a set $\mathcal{S}$ of dendric return morphisms. Recall that Theorem 43 ensures that all morphisms in $\mathcal{S}$ can be assume to be tame, i.e., are compositions of permutations and of the elementary morphisms
@@
Furthermore, there exists a regular language over the alphabet of elementary morphisms that exactly corresponds to the set $\mathcal{S}_{3}=\{\alpha,\beta,\gamma,\delta_{n},\zeta_{n},\eta\mid n\in\mathbb{N}\}$:

### v3

@@ 8.1. Extension graphs in codings of IET
A shift space $X$ over $\mathcal{A}$ is the natural coding of a regular interval exchange transformation with the pair of orders $\binom{\leq_{1}}{\leq_{2}}$ if and only if it is minimal, $\mathcal{A}\subseteq\mathcal{L}(X)$ and it satisfies the following conditions for every $w\in\mathcal{L}(X)$:
@@
This result can be reformulated as follows. Let $\leq^{L}$ and $\leq^{R}$ be two total orders on $\mathcal{A}$. A factor $w$ of a shift space $X\subseteq A^{\mathbb{Z}}$ is said to be planar for $(\leq^{L},\leq^{R})$ if item 1 is satisfied for $\binom{\leq^{R}}{\leq^{L}}$. This implies that, placing the left and right vertices of $\mathcal{E}_{X}(w)$ on parallel lines and ordering them respectively by $\leq^{L}$ and $\leq^{R}$, the edges of $\mathcal{E}_{X}(w)$ may be drawn as straight noncrossing segments, resulting in a planar graph. In particular, $\mathcal{E}_{X}(w)$ is acyclic.
@@
    Any long enough left special factor $w$ is such that $E^{L}_{X}(w)$ is equal to $\{a,b\}$, where $a,b$ are consecutive for $\leq^{L}$. Furthermore, for every two $\leq^{L}$-consecutive letters $a,b$ and for all $n\in\mathbb{N}$, there is a (unique) left special factor $w$ of length $n$ such that $\{a,b\}\subseteq E_{X}^{L}(w)$.
@@
    The same holds on the right, i.e., any long enough right special factor $w$ is such that $E^{R}_{X}(w)$ is equal to $\{a,b\}$, where $a,b$ are consecutive for $\leq^{R}$. Furthermore, for every two $\leq^{R}$-consecutive letters $a,b$ and for all $n\in\mathbb{N}$, there is a (unique) right special factor $w$ of length $n$ such that $\{a,b\}\subseteq E_{X}^{R}(w)$.
@@ 8.2. S-adic characterization of regular interval exchange transformations
From Theorem 63, we know that the natural coding of a regular interval exchange transformation is a minimal dendric shift. By Theorem 1, any of its $\mathfrak{S}$-adic representation made of return morphisms labels a path in $\mathcal{G}^{L}(\mathfrak{S})$ and in $\mathcal{G}^{R}(\mathfrak{S})$. In this section, we characterize those labeled paths.
@@
Let $a,b$ be two distinct letters in $\mathcal{A}$. Let $\mathfrak{S}$ be a set of dendric return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$. The graph $\mathcal{G}_{IET}(\mathfrak{S})$ is such that its vertices are the irreducible pairs of orders $(\leq^{L},\leq^{R})$ such that $a<^{L}b$, and there is an edge from $(\leq^{L},\leq^{R})$ to $(\preceq^{L},\preceq^{R})$ labeled by $\sigma\in\mathfrak{S}$ if
@@
    $(G(\leq^{L}),\sigma,G(\preceq^{L}))$ is an edge of $\mathcal{G}^{L}(\mathfrak{S})$;
@@
    $(G(\leq^{R}),\sigma,G(\preceq^{R}))$ is an edge of $\mathcal{G}^{R}(\mathfrak{S})$;
@@
Let $\mathfrak{S}$ be a family of dendric return morphisms from $\mathcal{A}^{*}$ to $\mathcal{A}^{*}$ and let $X$ be a shift space having an $\mathfrak{S}$-adic representation ${\boldsymbol{\sigma}}=(\sigma_{n})_{n\geq 0}$. Then $X$ is the coding of a regular interval exchange if and only if ${\boldsymbol{\sigma}}$ is primitive and labels an infinite path in the graph $\mathcal{G}_{IET}(\mathfrak{S})$. Moreover, if this path starts in the pair $(\leq^{L},\leq^{R})$, then $X$ is the coding of a regular interval exchange for the orders $\binom{\leq^{R}}{\leq^{L}}$.
@@
Let us assume that $X$ is the coding of a regular interval exchange for the orders $\binom{\leq^{R}_{0}}{\leq^{L}_{0}}$, we can moreover assume that $a<^{L}_{0}b$, where $a$ and $b$ are the letters used to define $\mathcal{G}_{IET}(\mathfrak{S})$. Then, $\sigma_{0}$ is $(\leq^{L}_{0},\leq^{R}_{0})$-planar, and by Lemma 67, there exist two orders $\leq^{L}_{1}$ and $\leq^{R}_{1}$ such that $a<^{L}_{1}b$ and $\sigma$ is planar preserving from $(\leq^{L}_{1},\leq^{R}_{1})$ to $(\leq^{L}_{0},\leq^{R}_{0})$ or from $((\leq^{L}_{1})^{*},(\leq^{R}_{1})^{*})$ to $(\leq^{L}_{0},\leq^{R}_{0})$.
As $X_{\boldsymbol{\sigma}}^{(1)}$ is dendric, by Corollary 70, it is the coding of a regular interval exchange for the orders $\binom{\leq^{R}_{1}}{\leq^{L}_{1}}$. In particular, the pair $(\leq^{L}_{1},\leq^{R}_{1})$ is irreducible. Moreover, $\sigma_{0}$ labels an edge from $G(\leq^{L}_{0})$ to $G(\leq^{L}_{1})$ in $\mathcal{G}^{L}(\mathfrak{S})$ and from $G(\leq^{R}_{0})$ to $G(\leq^{R}_{1})$ in $\mathcal{G}^{R}(\mathfrak{S})$. Therefore, $\sigma_{0}$ labels an edge from $(\leq^{L}_{0},\leq^{R}_{0})$ to $(\leq^{L}_{1},\leq^{R}_{1})$ in $\mathcal{G}_{IET}(\mathfrak{S})$. We then iterate this reasoning to show that ${\boldsymbol{\sigma}}$ labels an infinite path in $\mathcal{G}_{IET}(\mathfrak{S})$.
Assume now that ${\boldsymbol{\sigma}}$ is primitive and labels a path $((\leq^{L}_{n},\leq^{R}_{n}))_{n\geq 0}$ in $\mathcal{G}_{IET}(\mathfrak{S})$. Since ${\boldsymbol{\sigma}}$ also labels an infinite path in $\mathcal{G}^{L}(\mathfrak{S})$ and $\mathcal{G}^{R}(\mathfrak{S})$, $X$ is minimal dendric. Let us prove that it is planar for $(\leq^{L}_{0},\leq^{R}_{0})$. The $\sigma_{0}$-initial factors of $X$ are planar for $(\leq^{L}_{0},\leq^{R}_{0})$ since $\sigma_{0}$ is $(\leq^{L}_{0},\leq^{R}_{0})$-planar. Let $u\in\mathcal{L}(X)$ be a non-$\sigma_{0}$-initial factor of $X$. There exist a unique $k>0$ and a unique $v\in\mathcal{L}(X_{\boldsymbol{\sigma}}^{(k)})$ $\sigma_{k}$-initial such that $u$ is an extended image of $v$ by $\sigma_{0}\dots\sigma_{k-1}$ by Proposition 8. However, $\sigma_{k}$ is $(\leq^{L}_{k},\leq^{R}_{k})$-planar thus $v$ is planar for these orders. Iterating Proposition 69 $k$ times allows us to conclude that $u$ is planar for $(\leq^{L}_{0},\leq^{R}_{0})$. As this is true for any $u\in\mathcal{L}(X)$, $X$ is the coding of a regular interval exchange for the orders $\binom{\leq^{R}_{0}}{\leq^{L}_{0}}$. ∎
@@ 9. Future work
Given a set $\mathfrak{S}$ of dendric return morphisms, Theorem 1 characterizes the $\mathfrak{S}$-adic shifts that are minimal and dendric. We thus directly obtain the following corollary.
@@
Let $\mathfrak{S}$ be a set of dendric return morphisms defined on some alphabet $\mathcal{A}$ such that any minimal dendric shift over $\mathcal{A}$ admits an $\mathfrak{S}$-adic representation. Then a shift space $X\subseteq\mathcal{A}^{\mathbb{Z}}$ is minimal and dendric if and only if it has a primitive $\mathfrak{S}$-adic representation labeling paths in the graphs $\mathcal{G}^{L}(\mathfrak{S})$ and $\mathcal{G}^{R}(\mathfrak{S})$.
This corollary is however not completely satisfactory. Indeed it is in general difficult to find such a set $\mathfrak{S}$ of dendric return morphisms. Recall that Theorem 43 ensures that all morphisms in $\mathfrak{S}$ can be assume to be tame, i.e., are compositions of permutations and of the elementary morphisms
@@
Furthermore, there exists a regular language over the alphabet of elementary morphisms that exactly corresponds to the set $\mathfrak{S}_{3}=\{\alpha,\beta,\gamma,\delta_{n},\zeta_{n},\eta\mid n\in\mathbb{N}\}$: